### v1

# Evidence for Inflation

Abstract
Inflation as the leading paradigm depicting the very early universe physics could leave imprints on the cosmic microwave background (CMB) radiation. Using currently available CMB observations, we give the first direct evidence of inflation. We discuss the theoretical implications of our results including the energy scale of inflation, inflaton field excursion, Hubble expansion rate during inflation, equation of state of inflation, primordial tensor non-Gaussianity, primordial tensor power spectrum, B-mode anisotropy and inflationary gravitational wave background.
Introduction. The biggest question in science is the origin of the universe. In the hot big bang paradigm, shortly after the singularity, the universe is assumed to undergo an quasi-exponential expansion phase around $10^{-34}$ s, i.e., the so-called inflation Brout:1977ix ; Starobinsky:1980te ; Kazanas:1980tx ; Sato:1981qmu ; Guth:1980zm ; Linde:1981mu ; Linde:1983gd ; Albrecht:1982wi , which sets the very special initial conditions of such a huge universe we observe today. After inflation, the universe experiences in succession the (p)reheating, radiation-dominated, recombination, dark ages, reionization and dark energy eras. Clearly, as the onset of the universe, inflation will leave imprints on all the observations from a later stage. Although the nature of inflation is still unknown, we can probe the earliest moments of the universe now, due to well theoretical modeling and high precision observations Martin:2013tda ; Achucarro:2022qrl ; Baumann:2009ds . To decode the physics of inflation, the main tool is the statistics of primordial scalar and tensor fluctuations, where the former seeds the formation of large scale structure (LSS) and induces the primordial features Slosar:2019gvt ; Chluba:2015bqa and primordial non-Gaussianity (PNG) Maldacena:2002vr ; Meerburg:2019qqi ; Chen:2010xka ; Collins:2011mz , while the latter produces the primordial gravitational waves (PGWs) Achucarro:2022qrl ; Seljak:1996gy ; Krauss:2010ty ; Kamionkowski:2015yta ; Caprini:2018mtu as the earliest ripples of spacetime. These three important predictions of inflation help extract physical information at different inflationary energy scales and could be detected by the present-day CMB and LSS observations Krauss:2010ty . Up to now, the Planck CMB data Planck:2019nip ; Planck:2018lbu has constrained the deviation from the almost scale-invariant primordial scalar spectrum to be less than one percent of the amplitude $A_{s}$ Planck:2018vyg ; Planck:2018jri and shown no evidence of PNG Planck:2019kim . Very excitingly, using a logarithmic prior of the tensor-to-scalar ratio $r$, we have recently given the evidence of PGWs at beyond the $5\,\sigma$ confidence level (CL) via four available CMB experiments including the Planck Planck:2018vyg ; Planck:2018jri , WMAP WMAP:2012nax , Atacama Cosmology Telescope (ACT) ACT:2020frw ; ACT:2020gnv and the South Pole Telescope (SPT) SPT-3G:2021eoc ; SPT-3G:2022hvq , and provided the corresponding $1\,\sigma$ constraints on $\log_{10}r$ Wang:2024hhh , which can act as the direct evidence of inflation and have outstanding implications on the very early universe physics.
In this work, we present the new evidence of PGWs from the BICEP/Keck Array CMB B-mode polarization data BICEP:2021xfz and novel possibilities from our data analysis and, more importantly, interpret the theoretical implications of our constraining results of PGWs in the framework of the single-field slow-roll inflation, including the energy scale of inflation, inflaton field excursion, Hubble expansion rate during inflation, inflationary equation of state (EoS), primordial tensor non-Gaussianity, primordial tensor power spectrum (PTPS), angular power spectrum of CMB B-modes and inflationary gravitational wave background (IGWB).
New evidences and novel possibilities. Following the analysis methodology of the Planck collaboration Planck:2018vyg ; Planck:2018jri and our previous work Wang:2024hhh , we write the PTPS as
$\mathcal{P}_{T}(k)=A_{t}\left(\frac{k}{k_{p}}\right)^{n_{t}+\frac{1}{2}n_{trun}\ln\left(\frac{k}{k_{p}}\right)},$ (1)
where $k$, $k_{p}$, $A_{t}$, $n_{t}$ and $n_{trun}$ are, respectively, the comoving wavenumber, tensor pivot scale, amplitude of PTPS, tensor spectral index and running of tensor spectral index. $r\equiv A_{t}/A_{s}$, where $A_{s}$ denotes the amplitude of primordial scalar power spectrum. The next-order inflation consistency relation (ICR) employed by the Planck collaboration, namely $n_{t}=-r(2-r/8-n_{s})/8$ and $n_{trun}=r(r/8+n_{s}-1)/8$ Planck:2018vyg ; Planck:2018jri ; Planck:2015sxf is used in this work, where $n_{s}$ is the scalar spectral index.
Figure 1: Left. The two-dimensional marginalized posterior distributions of the parameter pair ($n_{s}$, $\log_{10}r$) from the Planck and BK18 observations. We consider the cases of next-order ICR and $n_{t}=0$ when using the Planck data. Right. The derived constraint on the tensor spectral index $n_{t}$ from the Planck data for a single-field slow-roll inflation with the next-leading-order ICR.
Figure 2: The schematic plot for the relations between a single-field slow-roll inflation with the next-order ICR (red) or with next-order ICR but $n_{t}=0$ (orange) and primordial tensor features including $\Lambda$CDM+$r$+$n_{t}$, $\Lambda$CDM+$r$+$n_{t}$+$n_{trun}$ and extended tensor features.
To implement the Bayesian analysis, we adopt the following two datasets: (i) Planck. The Planck 2018 $-\ell$ plik temperature (TT) data at multipoles $30\leqslant\ell\leqslant 2508$, polarization (EE) and their cross-correlation (TE) likelihoods at $30\leqslant\ell\leqslant 1996$, and the $-\ell$ TT Commander and SimAll EE data at $2\leqslant\ell\leqslant 29$ Planck:2019nip as well as the Planck lensing likelihood Planck:2018lbu from the SMICA maps at $8\leqslant\ell\leqslant 400$. (ii) BK18. The BICEP2, Keck Array and BICEP3 CMB polarization data up to the 2018 observing season BICEP:2021xfz . This dataset aiming at detecting the CMB B-modes contains the additional Keck Array observations at 220 GHz and BICEP3 observations at 95 GHz relative to the previous 95/150/220 GHz data.
Furthermore, we take the public Boltzmann code CAMB Lewis:1999bs to compute the theoretically cosmological quantities and adopt the Monte Carlo Markov Chain (MCMC) method to infer the posterior probability density distributions of model parameters using the package CosmoMC Lewis:2002ah ; Lewis:2013hha . The generated MCMC chains is post-processed by the package Getdist Lewis:2019xzd and the corresponding convergence rule of each MCMC run is the Gelman-Rubin diagnostic $R-1\lesssim 0.02$ Gelman:1992zz . Same as our previous analysis Wang:2024hhh , the following uniform priors for seven model parameters are used: the the logarithmic tensor-to-scalar ratio $\log_{10}r\in[-30,5]$, $\mathrm{ln}(10^{10}A_{s})\in[2,4]$, $n_{s}\in[0.8,1.2]$, baryon fraction $\Omega_{b}h^{2}\in[0.005,0.1]$, cold dark matter fraction $\Omega_{c}h^{2}\in[0.001,0.99]$, comoving acoustic angular scale at the recombination epoch $100\theta_{MC}\in[0.5,10]$ and optical depth due to reionization $\tau\in[0.01,0.8]$. The pivot scale used is 0.05 Mpc-1.
Figure 3: The constrained primordial tensor power spectrum (left) and CMB B-mode angular power spectrum (right) with $1\,\sigma$, $2\,\sigma$ and $3\,\sigma$ confidence regions. The magenta and green shaded regions denote the CMB scales and the cosmic variance, respectively. The red solid and blue dashed lines represent the best fit and the pivot scale $k_{p}=0.05$ Mpc-1, respectively.
Our previous results Wang:2024hhh reveals that PGWs have been detected by four CMB experiments including Planck, WMAP, ACT and SPT separately as well as their combinations with BK18 and DESI observations DESI:2024mwx ; DESI:2024uvr ; DESI:2024lzq at beyond the $5\,\sigma$ CL. However, the constraint on six free parameters of $\Lambda$CDM plus $\log_{10}r$ from BK18 alone is not investigated, even though it can not give any restriction on the scalar spectral index. Instead, we just constrain $\log_{10}r$ by assuming the Planck 2018 fiducial cosmology. In this study, we show the BK18 constraint on $\Lambda$CDM+$\log_{10}r$ in Fig.1, which gives $\log_{10}r=-6.9^{+5.0}_{-2.7}$ at the $1\,\sigma$ CL and $\log_{10}r=-6.9^{+5.6}_{-6.0}$ at the $2\,\sigma$ CL. Actually, same as our previous results Wang:2024hhh , BK18 also give the detection of PGWs at beyond the $5\,\sigma$ CL (see the supplementary material) but with smaller lower bounds than the Planck constraint. It is noteworthy that all the above results are obtained by assuming the second-order ICR in a single-field slow-roll inflation. Interestingly, there is a novel possibility that PGWs can still exist when fixing $n_{t}=0$. The reason why we consider this case is that current cosmological data only provides a very weak constraint on free $n_{t}$, which does not rule out $n_{t}=0$ Planck:2019kim . Although $n_{t}=0$ is different from the theoretical prediction of inflation $n_{t}<0$, it has interesting physical properties, namely reducing the primordial tensor parameter space and only introducing high order scale dependence of PTPS. In light of the Planck data, we find the $1\,\sigma$ and $2\,\sigma$ constraints $\log_{10}r=-4.1^{+1.4}_{-2.0}$ and $\log_{10}r=-4.1^{+2.6}_{-2.5}$ for the case of $n_{t}=0$, respectively. It is clear to see that this constraint gives a larger lower bound on $\log_{10}r$ than the case of full ICR by $\sim 3$ orders of magnitude. As expected, the primordial tensor parameter space or the posterior of $r$ is, to a large extent, compressed, implying that such a tensor feature may be easier to be detected or falsified by future high-sensitivity experiments.
As mentioned above, one of the most important prediction is a negative $n_{t}$. Consequently, we study this point in a single-field slow-roll inflation with next-order ICR using the Planck observations, and derive the constraint ${\rm log}_{10}(-n_{t})=-6.4\pm 2.5$ at the $1\,\sigma$ CL and ${\rm log}_{10}(-n_{t})=-6.4\pm 4.1$ at the $2\,\sigma$ CL. In Fig.1, one can easily find that $n_{t}<0$ is well consistent with the prediction of inflation, even if the value of $n_{t}$ is very small and close to zero. Therefore, $r>0$ and $n_{t}<0$ obtained here at beyond the $5\,\sigma$ CL can act as the first direct evidence of inflation.
So far, the detection of PGWs depends strongly on the next-order ICR regardless of whether $n_{t}=0$ or not. But, we do not explore whether PGWs also exist in extended parameter spaces such as $\Lambda$CDM+$r$+$n_{t}$, $\Lambda$CDM+$r$+$n_{t}$+$n_{trun}$ or more tensor features. We show roughly the relations between models considered here and possibly extended features in Fig.2. The reason why there is an overlap between $r$+ICR and $r$+$n_{t}$ is that the posterior of $n_{t}$ in $r$+ICR is a small subset of that in $r$+$n_{t}$ and the posterior of $n_{trun}$ in $r$+ICR includes $n_{trun}=0$ in $r$+$n_{t}$. Naturally, $r$+ICR with $n_{t}=0$ and $r$+$n_{t}$ display a similar overlap. The only difference is that $n_{t}$ lives in a small negative range satisfying the theoretical prediction of inflation in $r$+ICR relative to $r$+ICR with $n_{t}=0$. This is the reason why they have no any overlap. Since PGWs have been detected with a high statistical significance by five independent CMB facilities Planck:2018vyg ; Planck:2018jri ; WMAP:2012nax ; ACT:2020frw ; ACT:2020gnv ; SPT-3G:2021eoc ; SPT-3G:2022hvq ; BICEP:2021xfz , it is very necessary to probe completely the primordial tensor parameter space using the same methodology and datasets depicted here. An in-depth exploration of whether there are signals of primordial tensor features is forthcoming Wang:2024kkk .
Energy scale. The detection of PGWs can provide a unique way to estimate the energy scale of inflation, which is expressed as $V_{p}=3\pi^{2}A_{s}/2rM_{\rm pl}^{4}=1.01\times 10^{16}{\rm GeV}(r/0.01)^{1/4}$ Planck:2018jri ($M_{\rm pl}$ is the reduced Planck mass) at the horizon exit of a pivot scale, since they carry the information of their generating mechanism. Using the constraints on $\log_{10}r$ from Planck Wang:2024hhh , we obtain the inflationary energy scale as $V_{p}^{1/4}=1.35^{+4.36}_{-1.03}\times 10^{15}$ GeV at the $1\,\sigma$ CL and $V_{p}^{1/4}=1.35^{+12.95}_{-1.22}\times 10^{15}$ GeV at the $1\,\sigma$ CL, respectively. Planck’s $1\,\sigma$ constraint indicates that the energy scale of inflation is lower than the Grand Unified Theory (GUT) scale $\sim 10^{16}$ GeV. Interestingly, its $2\,\sigma$ upper limit translated to $1.43\times 10^{16}$ GeV is higher than the GUT scale. This implies that the electromagnetic, weak and strong forces could be unified during inflation. Although the energy scale is lower than the Planck scale $\sim 10^{19}$ GeV, it is far larger than $\sim 10^{4}$ GeV that can be reached by the most powerful high-energy particle facility on earth, i.e., the Large Hadron Collider (LHC). More intriguingly, our findings reveals that the physics before inflation should be more close to the Planck scale.
Hubble expansion rate. Using the above constraints on the energy scale of inflation and the Friedman equation $H_{p}^{2}\approx V_{p}/(3M_{\rm pl}^{2})$, the Hubble parameter during inflation can be constrained to be $H_{p}/M_{\rm pl}=1.78^{+29.90}_{-1.68}\times 10^{-7}$ at the $1\,\sigma$ CL and $H_{p}/M_{\rm pl}=1.78^{+198.22}_{-1.76}\times 10^{-7}$ at the $2\,\sigma$ CL. The $2\,\sigma$ upper bound of $2.0\times 10^{-5}$ is tighter than $2.5\times 10^{-5}$ obtained by the Planck collaboration Planck:2018jri .
Figure 4: The constrained inflationary gravitational wave background with $1\,\sigma$, $2\,\sigma$ and $3\,\sigma$ confidence regions. The magenta shaded region, red solid line and blue dashed line denote the CMB scales and the best fit and the pivot scale $k_{p}=0.05$ Mpc-1, respectively.
Inflaton excursion. The scalar field excursion $\Delta\phi$ characterizing the variation of inflaton field between the horizon exit of the pivot scale and the end of inflation provides the information about the viable quantum gravity theory. It is related to the tensor-to-scalar ratio via $\Delta\phi/M_{\rm pl}=(r/8)^{1/2}N_{e}$ Lyth:1996im , where $N_{e}$ is the e-folding number during this period. The constraint of this field net excursion leads to a classification of larger-field and small-field inflationary models, where the discriminating value is the Planck mass. Choosing $N_{e}=30$ Boubekeur:2005zm ; Boubekeur:2012xn , Planck data gives the $1\,\sigma$ and $2\,\sigma$ constraints $\Delta\phi/M_{\rm pl}=0.019^{+0.316}_{-0.018}$ and $0.019^{+2.097}_{-0.019}$, respectively. One can easily find that the result is consistent with effective field theories supporting small-field excursions at the $1\,\sigma$ CL after removing the higher-order corrections by introducing some symmetries, and that the $2\,\sigma$ upper bound of $\Delta\phi/M_{\rm pl}=2.116$ from Planck is higher than the Planck scale implying a possible existence of super-Planckian field excursion, which poses a great challenge for theoretical modeling. Interestingly, our constraint is compatible with the first criterion of the swampland conjecture (SC) Ooguri:2006in ; Obied:2018sgi ; Agrawal:2018own , i.e., $|\Delta\phi/M_{\rm pl}|<\Delta\sim\mathcal{O}(1)$ with $\Delta$ being a constant at the $2\,\sigma$ CL. However, it can place a strong constraint on the constant $c$ of the second criterion of SC, $M_{\rm pl}V^{\prime}/V>c\sim\mathcal{O}(1)$, and gives $c<5.60\times 10^{-6}\ll 1$ at $2\,\sigma$ level using the relation $r>8c^{2}$ Matsui:2018bsy ; Ben-Dayan:2018mhe . Clearly, this bound poses a very strong tension with SC indicating that string-based inflationary models are likely not viable Wang:2018duq .
Nature of inflation. The tensor spectral index can act as a function of inflation EoS $\omega$ via the formula $n_{t}=4/(1+3\omega)+2$ Mukhanov:1990me ; Lyth:1998xn . Using the Planck data, we obtain the $1\,\sigma$ and $2\,\sigma$ constraint ${\rm log}_{10}(1+\omega)=-6.9\pm 2.5$ and $-6.9\pm 4.1$ (see the supplementary material) implying that $w>-1$ but is very close to $\omega=-1$. This means that the nature of the single-field slow-roll inflation with the next-order ICR is quintessence. Note that we only study the phenomenological nature of inflation here and do not consider its fundamental physics origin. As demonstrated in Wang:2024hhh , $R^{2}$ inflation with a large e-folding number, say $N=60$, is well supported by observations. Interestingly, besides $R^{2}$ inflation, there is still a large part of parameter space of $(n_{s},\log_{10}r)$ needed to be explained by new physics.
Non-Gaussianity. The theory predicts that primordial scalar and tensor non-Gaussianities have the forms of $f_{\rm NL}^{s}=5/12(n_{s}-1)$ and $f_{\rm NL}^{t}=5/24(3-n_{t})$ after discarding the higher order terms Copeland:1994vg ; Dine:1995uk ; Maldacena:2002vr ; Kundu:2013gha . The Planck data has constrained $n_{s}$ to a $0.4\%$ precision and gives $f_{\rm NL}^{s}=-0.0146\pm 0.0018$ at $1\,\sigma$ level Planck:2019kim . Note that this constraint is obtained without inflationary model dependence. However, the tensor non-Gaussianity $f_{\rm NL}^{t}=0.6250^{+2.6144\times 10^{-5}}_{-8.2677\times 10^{-8}}$ derived in a single-field slow-roll inflation has been constrained to a very high precision at the $1\,\sigma$ CL.
PTPS. Since our results provide a clear lower bound on the PTPS amplitude, we can give the observationally allowed PTPS across different scales. In Fig.3, we present the constrained PTPS within the $3\,\sigma$ CL from Plank by using the best fit $n_{t}=-10^{-6.4}$ and $n_{trun}=-10^{-4.85}$ derived from the ICR. One can easily find that PTPS is scale-invariant due to very small $n_{t}$ and $n_{trun}$ and spans almost ten orders of magnitude at $3\,\sigma$ level. This will lead to many underlying possibilities of the very early universe to interpret the observed PTPS.
B-mode anisotropy. Similar to PTPS, in Fig.3, we give the allowed angular power spectrum (APS) of CMB B-mode by Planck at $3\,\sigma$ level. Since B-mode mainly contributes to the CMB temperature APS at large scales, we study the B-mode APS at $2002<\ell<$ 200. The constrained APS spans about three orders of magnitude within $3\,\sigma$ CL at the largest scale, say $\ell=2$, however, this large uncertainty span reduces to zero around $\ell=200$. We also show the anisotropic B-mode maps for the best fit, $1\,\sigma$, $2\,\sigma$ and $3\,\sigma$ upper bounds of $\log_{10}r$ from the Planck-only constraint in the supplementary material.
IGWB. The stochastic gravitational wave background (SGWB) can be probed by pulsar timing arrays (PTA) Verbiest:2021kmt and gravitational wave observatories LIGOScientific:2016jlg . Recently, several PTA experiments such as NANOGrav NANOGrav:2023gor , EPTA Antoniadis:2023rey , PPTA Reardon:2023gzh and CPTA Xu:2023wog have reported the substantial evidences of SGWB. Interestingly, inflation can act as a possible cosmological source of SGWB. Although IGWB as a single source has been constrained by PTA data NANOGrav:2023hvm , we argue that IGWB may not be the only source of SGWB. Specifically, we employ the IPTA DR2 Antoniadis:2022pcn to constrain the single-field slow-roll inflation with next-order ICR and find ${\rm{log}_{10}}r=0.485^{+0.013}_{-0.010}$ at $1\,\sigma$ level (see the supplementary material), which is much larger than the constraint from CMB data. Hence, we conclude that IGWB really exist in the universe but must be mixed with other astrophysical or cosmological sources to explain the SGWB. Moreover, considering appropriately the contribution of radiation-matter equality and reheating eras to the transfer function Guzzetti:2016mkm , the gravitational wave energy spectrum constrained by Planck data is shown in Fig.4. Interestingly, the energy densities of gravitational waves $\Omega_{\rm gw}h^{2}$ span $\sim 8$ orders of magnitudes over all the frequencies at $3\,\sigma$ level. Especially, when frequencies are larger than CMB ones, $\Omega_{\rm gw}h^{2}$ shows a clear frequency-invariant feature.
Concluding remarks. The detection of PGWs plays a key role in understanding the connection between general relativity and quantum theory and probing the high energy physics close to the Planck scale. It can provide the direct evidence of the leading paradigm of the very early universe, i.e., inflation. By presenting new evidences and novel possibilities of PGWs, we increases the correctness of inflation that really exists at the beginning of the universe in light of the fact that four CMB experiments have detected the PGWs separately. However, we can not conclude that inflationary mechanism has been completely confirmed by current observations, since primordial scalar and tensor non-Gaussianities as the important prediction in a single-field slow-roll inflation are not yet detected.
Furthermore, current data can not give a good constraint on the extended features such as $\Lambda$CDM+$n_{t}$+$n_{trun}$, and consequently, there is still a large tensor parameter space need to be explored by future observations and theoretical modeling. In the above, we just mention that $R^{2}$ inflation can explain the current constraints well. Actually, we require a large number of models to be proposed to interpret the large $n_{s}$-$\log_{10}r$ contour, where $r$ spans several orders of magnitude.
Since our constraining results have revealed that the GUT scale can be reached during inflation within $2\,\sigma$ level, the particle production of dark matter around the GUT scale is allowed by observations. As a consequence, our constraints could inspire the exploration of dark matter model building at extremely high energy scales, and intriguingly, provide a new window to investigate the physics before inflation.
Acknowledgements. DW is supported by the CDEIGENT Fellowship of Consejo Superior de Investigaciones Científicas (CSIC). DW warmly thanks the usage of Planck Planck:2018vyg , WMAP WMAP:2012nax , ACT ACT:2020gnv , SPT SPT-3G:2022hvq , BK18 BICEP:2021xfz and DESI DESI:2024mwx data.
S1 Supplementary Material
S1.1 A. $5\,\sigma$ contours from BK18
In Fig.S1, we present the constrained ($n_{s}$, $\log_{10}r$) parameter space from the BK18 observations within the $5\,\sigma$ confidence level (CL). It is easy to see that PGWs are detected at beyond the $5\,\sigma$ CL. This can act as an important evidence of PGWs independent of the four CMB experiments including Planck, WMAP, ACT and SPT.
S1.2 B. Inflationary equation of state
In Fig.S2, we show the one-dimensional posterior distribution of inflationary equation of state (EoS). It is easy to see that the nature of inflation is quintessence, i.e., $\omega>-1$, at beyond the $5\,\sigma$ CL.
S1.3 C. CMB B-mode polarization maps
In Fig.S3, we show the CMB B-mode polarization maps for the best fit, $1\,\sigma$, $2\,\sigma$ and $3\,\sigma$ upper bounds of $\log_{10}r$ from the Planck-only constraint.
Figure S1: The two-dimensional marginalized posterior distributions of the parameter pair ($n_{s}$, $\log_{10}r$) from the BK18 observations within the $5\sigma$ confidence level.
Figure S2: The one-dimensional marginalized posterior distribution of inflationary equation of state from the Planck data.
Figure S3: The CMB B-mode polarization maps for the best fit (upper left), $1\,\sigma$ (upper right), $2\,\sigma$ (lower left) and $3\,\sigma$ (lower right) upper bounds of $\log_{10}r$ from the Planck-only constraint.
Figure S4: The one-dimensional marginalized posterior distribution of the logarithmic tensor-to-scalar ratio ${\rm{log}_{10}}r$ from the IPTA DR2.
S1.4 D. The inconsistency of constraints on $r$ between CMB and pulsar timing arrays
Using the formula of the gravitational wave energy spectrum shown in [] and employing the analysis methods used in [], we give the constraint on the tensor-to-scalar ratio ${\rm{log}_{10}}r=0.485^{+0.013}_{-0.010}$ at the $1\,\sigma$ CL, ${\rm{log}_{10}}r=0.485^{+0.023}_{-0.028}$ at the $2\,\sigma$ CL and ${\rm{log}_{10}}r=0.485^{+0.027}_{-0.031}$ at the $3\,\sigma$ CL via the international pulsar timing array data release two (IPTA DR2) []. One can easily find that they are far larger than the CMB constraints and consequently, PTA is in a very strong tension with CMB when constraining the single-field slow-roll inflation.